Circular No. 8092

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2003 CP_{7} (LINEAR-NEAT) A comet discovered on NEAT Palomar images obtained on Mar. 10,

and posted on the NEO Confirmation Page, was reported by K. Lawrence as showing a nuclear condensation of diameter ∼ 7′′ and a tail ∼ 8′′ long toward the west. The cometary nature was confirmed by J. Young at Table Mountain on Mar. 12.4 UT. The Minor Planet Center has linked this object to an apparently asteroidal LINEAR object of mag 19.0 on Feb. 1 and 4 that was designated 2003 CP_{7} (MPS 73383–73384, Feb. 16).

2003 UT α_{2000} δ_{2000} m_{1} Observer
Feb. 1.39231 11^{h}27^{m}24.^{s}65 +11^{o}10′53′′.4  LINEAR
Mar. 10.35520 11 10 16.06 +15 50 08.0 17.7 NEAT

The available astrometry, the following elliptical orbital elements, and an ephemeris appear on MPEC 2003-E47.

T = 2003 Apr. 29.0579 TT  ω = 42.^{o}5367
e = 0.248745  Ω = 133.1425 }2000.0
q = 3.018060 AU  i = 12.3472
a = 4.017357 AU n^{o} = 0.1224034 P = 8.052 years

COMET C/2003 E1 (NEAT) S. Pravdo reports another NEAT comet discovery, the object having a

tail extending ∼ 8′′ in p.a. 215^{o} on Mar. 11. Young also found it cometary on Mar. 12. The available astrometry, very uncertain parabolic orbital elements [T = 2004 Mar. 13.3 TT, ω = 110^{o}.8, Ω = 141^{o}.9, i = 37^{o}.6 (equinox 2000.0), q = 2.950 AU], and an ephemeris appear on MPEC 2003-E48.

2003 UT  α_{2000}  δ_{2000}  m_{1}
Mar. 9.50577  12^{h}25^{m}51.^{s}32  +27^{o}07′41′′.1  19.7

SUPERNOVA 2003bu IN NGC 5393 Further to IAUC 8091, H. Pugh and W. Li report the LOTOSS dis-

covery of an apparent supernova on unfiltered KAIT images taken on Mar. 11.5 (mag ∼ 17.5) and 12.5 UT (mag ∼ 17.2). SN 2003bu is located at α = 14^{h}00^{m}31^{s}.45, δ = −28^{o}52′32′′.7 (equinox 2000.0), which is 7′′.3 west and 3′′.7 south of the nucleus of NGC 5393. A KAIT image taken on Feb. 10.5 showed nothing at this position (limiting mag ∼ 18.0).

2003 March 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green